A photon with energy 2.28 eV is absorbed by a hydrogen atom. Find (a) the minimum n for a hydrogen atom that can be ionized by such a photon and (b) the speed of the electron released from the state in part.
1
Expert's answer
2020-10-19T13:20:40-0400

A) since the energy of level "n" in the stationary state of hydrogen atom is defined as "E_n=\frac{m\times e^4}{32\times \pi^2\times\varepsilon_0^2\times \hbar^2 } \times \frac {1}{n^2}" , where "m" is the electron mass, "e" the electron charge, "\varepsilon_0" – electric constant, "\hbar" Planck's constant Dirac.

Express "n" and you get "n=\sqrt{\frac{m\times e^4}{32\times \pi^2\times\varepsilon_0^2\times h^2 } \times \frac {1}{ E_n }}=" "\sqrt{\frac{9.1\times 10^{-31}\times (1.6\times 10^{-19})^4}{32\times \pi^2\times(8.85\times10^{-12})^2\times (1.05\times 10{-34})^2 } \times \frac {1}{ 2.28\times 1.6\times 10^{-19}}}\approx2" .

B) Since the kinetic energy is "E=\frac{m\times v^2}{2}" ,hence "v=\sqrt{\frac{2\times 2.28\times 1.6\times 10^{-19}}{9.1\times 10^{-31}}}=8.95\times 10^5" m/s.